OpenStudy (anonymous):

Solve. 2x^2 + 5x = 12

OpenStudy (anonymous):

2x^2+5x-12=0 (2x-3)(X+4)=0\ X=-4, X=3/2

OpenStudy (anonymous):

Thank you again :)